OpenStudy (anonymous):

Write the equation 2x - 3y = 6 in slope-intercept form. y = 3/2x + 3 y = 2/3x + 3 y = 2/3x - 2

OpenStudy (anonymous):

first step add -2x to both sides and you'll get \[\LARGE 2x-2x-3y=6-2x\] now what can you simplify there?

OpenStudy (anonymous):

would it be the second one?

OpenStudy (anonymous):

not actually... try again

OpenStudy (anonymous):

you have -3y=-2x+6 divide both sides by -3 and you'll get: \[\LARGE \frac{-3y}{-3}=\frac{-2x}{-3}-\frac{6}{3}\]

OpenStudy (anonymous):

then...

OpenStudy (anonymous):

its y=2/3x-2

OpenStudy (anonymous):

oh yes it is.. I'm sorry I wasn't here.. :(